# Verifiably Following Complex Robot Instructions with Foundation Models

### v2

@@ Abstract
Enabling mobile robots to follow complex natural language instructions is an important yet challenging problem. People want to flexibly express constraints, refer to arbitrary landmarks and verify behavior when instructing robots. Conversely, robots must disambiguate human instructions into specifications and ground instruction referents in the real world. We propose Language Instruction grounding for Motion Planning (LIMP), an approach that enables robots to verifiably follow expressive and complex open-ended instructions in real-world environments without prebuilt semantic maps. LIMP constructs a symbolic instruction representation that reveals the robot’s alignment with an instructor’s intended motives and affords the synthesis of robot behaviors that are correct-by-construction. We perform a large scale evaluation and demonstrate our approach on 150 instructions in five real-world environments showing the generality of our approach and the ease of deployment in novel unstructured domains. In our experiments, LIMP performs comparably with state-of-the-art LLM task planners and LLM code-writing planners on standard open vocabulary tasks and additionally achieves 79% success rate on complex spatiotemporal instructions while LLM and Code-writing planners both achieve 38%. See supplementary materials and demo videos at robotlimp.github.io.
Keywords: Foundation Models, Mobile Manipulation, Task and Motion Planning
1 Introduction
Robots need a rich understanding of natural language to be usable by non-experts in unstructured environments. People, on the other hand, need to be able to verify that a robot has understood a given instruction and will act appropriately. However, natural language instructions can be quite complex, often featuring ambiguous phrasing, intricate spatiotemporal constraints, and unique referents. Consider the instruction: “Go to the building with windows while avoiding the red picnic tables, then bring me the book between the pine tree and the graffiti wall”, to solve this task a robot must ground open-vocabulary referents (“building with windows”, “red picnic tables”, “book”, “pine tree”, “graffiti wall”), adhere to temporal constraints (“then bring”, “while avoiding”), and disambiguate multiple instances of an object category using descriptive spatial information (“book between the pine tree and the graffiti wall”). The rise of foundation models in robotics [1, 2] offers a path to achieving such long-horizon goals, and indeed some approaches for robot instruction following have been proposed, but have largely been focused on navigation [3, 4, 5, 6, 7]. These methods, broadly classified under object goal navigation [8], enable navigation to instances of an object category, but are limited in localizing spatial references and disambiguating object instances based on descriptive language. Other works [9, 10, 11] that extend instruction following to mobile manipulation have been limited to tasks with simple temporal constraints expressed in unambiguous language. Existing efforts typically rely on Large Language Models (LLMs) as complete planners, bypassing intermediate symbolic representations that may provide verification of correctness prior to task execution. Alternative approaches that leverage code-writing LLMs [5, 6, 12] are susceptible to errors in generated code which may lead to unsafe/unaligned robot behaviors. Borrowing ideas from formal methods to map between natural language and specification languages like temporal logics [13] provides a robust framework for language disambiguation, handling complex temporal constraints, and behavior verification. However, prior works along this line require prebuilt semantic maps with discrete sets of prespecified referents/landmarks from which instructions can be constructed [7, 14, 15].
We propose Language Instruction grounding for Motion Planning (LIMP), a method that leverages foundation models and temporal logics to dynamically generate instruction-conditioned semantic maps that enable robots to construct verifiable controllers for following navigation and mobile manipulation instructions with open vocabulary referents and complex spatiotemporal constraints. In a novel environment, LIMP first constructs a 3D representation of the environment via SLAM, then leverages large language models to translate complex natural language instructions into temporal logic specifications with a novel composable syntax that enables referent disambiguation. Instruction referents are then detected and grounded using Visual Language Models (VLMs) and spatial reasoning. Finally, a task and motion plan is synthesized to progressively guide the robot through the subgoals required to satisfy the instruction, as shown in Figure LABEL:fig:splash. In summary, we make the following contributions: (1) A modular framework for translating expressive natural language instructions into temporal logic, grounding instruction referents, and executing the command via task and motion planning (TAMP). (2) A spatial grounding method for detecting and localizing open vocabulary objects with spatial constraints in 3D metric maps. (3) A TAMP algorithm that localizes regions of interest (goal/avoidance zones) and synthesizes constraint-satisfying motion plans to progressively solve long-horizon tasks.
2 Related Works
In this section, we briefly highlight the most relevant comparable works in the space of visual scene understanding, instruction following from natural language, and task and motion planning. We offer a more detailed literature review in our appendix A2.
Visual Scene Understanding: NLMap [10] presents a scene representation that grounds open-vocabulary language queries to spatial locations via pre-trained VLMs. While NLMap is suitable for handling complex descriptions of individual objects, it is fundamentally unable to handle instructions involving complex constraints between multiple objects since it does not capture object relationships. LIMP handles these more complicated instructions by using a novel spatial grounding module to incorporate spatial relationships between objects, and leverages task and motion planners to satisfy these constraints.
Natural Language Instruction Following: Lang2LTL [16] is a multi-stage LLM-based approach that performs entity extraction and replacement to translate natural language instructions into a temporal logic expression. However, Lang2LTL relies on a prebuilt semantic map that grounds expression symbols, whereas our approach uses a novel scene representation that supports open-vocabulary language and generates the relevant landmarks based on the open-vocabulary instruction.
Task and Motion Planning: Action-Oriented Semantic Maps (AOSMs) [17] are a class of semantic maps that include additional models in space for where the robot can execute manipulation skills (represented as symbols), and are designed to work in conjunction with TAMP solvers for mobile manipulation domains. Our scene representation is similar to an AOSM since it captures spatial information about semantic objects of interest, and is compatible with TAMP solvers. However, once an AOSM is generated for a scene, there is only a closed-set of goals that can be planned for, whereas our approach can handle open-vocabulary task specifications.
3 Background
LIMP translates natural language instructions into temporal logic specifications for verifiable task and motion planning. Our approach is compatible with various specification languages and planning frameworks, however we choose Linear Temporal Logic (LTL) [18] to represent instructions due to its proven expressivity in representing a wide range of complex robot mission requirements [19]. LTL offers a rich and expressive syntax for defining the temporal properties of systems. Specifications consist of atomic propositions that reflect an aspect of a system’s state and the logical operators: negation ($\neg$), conjunction ($\land$), disjunction ($\lor$), implication ($\rightarrow$), in addition to temporal operators: next ($\mathcal{X}$), until ($\mathcal{U}$), globally ($\mathcal{G}$) and finally ($\mathcal{F}$). Despite its expressivity, LTL has seen limited use due to the expert knowledge required to construct specifications, however recent works have seen significant success removing this barrier by directly translating natural language into LTL [7, 20, 14, 21, 22, 23].
Behavior Verification: Expressing instructions as temporal logic specifications allows us to verify the correctness of generated plans a priori. However, instead of secondary explicit verification via approaches such as model checking, we follow conventional wisdom [24] and directly use specifications to synthesize plans that are correct-by-construction [25, 26].
4 Problem Definition
Given a natural language instruction $l$, our goal is to synthesize and sequence navigation and manipulation behaviors to produce a policy that satisfies the temporal and spatial constraints expressed in $l$. Spatial constraints determine the success of the task based on the sequence of poses the robot traversed during task execution, and temporal constraints determine the sequencing of spatial constraints as a function of task progression. We assume that a robot equipped with an RGBD camera has navigated around some space and captured a collection of images along with camera poses, with this data we build a metric map $m$ (e.g: pointcloud, 3D voxelgrid) of the environment. This map defines the space of possible $SE(3)$ poses $P$ and can be used to localize the robot (i.e: estimate $p_{robot}\in P$). Noteably, unlike previous work that leverage temporal logics [16], we do not assume access to a semantic map with the locations of objects or predicates prespecified. Instead, we leverage two types of foundation models: a task-agnostic visual language model $\sigma$ that given an image and text provides a bounding box/segmentation of the image based on the text, and an auto-regressive large language model $\psi$ that samples likely language tokens based on a history of tokens.
Navigation: We formalize navigation as object-goal oriented continuous path planning problems, where the goal is to generate paths to a goal pose set $P_{goals}\subset P$ while staying in feasible regions ($P_{feasible}\subset P$) and avoiding infeasible regions ($P_{infeasible}=P_{feasible}^{C}$). Infeasible regions include environment obstacles as well as dynamically determined semantic regions that violate constraints in the instruction $l$.
Manipulation: We formalize interactive manipulation behaviours as options [27] parameterized by objects. Consider an object parameter $\theta$ that parameterizes an option $O_{\theta}=(I_{\theta},\pi_{\theta},\beta_{\theta})$, the initiation set, policy, and termination condition are functions of both the robot pose $P$ and $\theta$. The initiation set $I_{\theta}$ denotes the global reference frame robot positions and object-centric attributes––such as object size––that determine if the option policy $\pi_{\theta}$ can be executed on the object $\theta$. To execute a manipulation skill on an object, an object-goal navigation behavior must first be synthesized and executed to bring the robot into proximity with the object in question. Our formalism is capable of encoding a wide variety of long-horizon object-centric tasks, but for simplicity, we demonstrate our approach with two types of tasks: multi-object goal navigation and open vocabulary mobile pick-and-place [9, 28]. These tasks require navigation, picking, and placing robot behaviors.
5 Language Instruction Grounding for Motion Planning
Figure 1: [A] LIMP translates natural language instructions into temporal logic expressions, where open-vocabulary referents are applied to predicates that correspond to robot skills––note the context-aware resolution of the phrase “blue one” to the referent “blue_sofa”. [B] Visual language models detect referents, while spatial reasoning disambiguates referent instances to generate a 3D semantic map that localizes instruction-specific referents. [C] Finally, the temporal logic expression is compiled into a finite-state machine, which a bi-level task and motion planner uses with dynamically-generated task progression semantic maps to progressively identify goals and constraints in the environment, and generate a plan that satisfies the high-level task specification.
LIMP interprets expressive natural language instructions and generates instruction-conditioned semantic maps that enable robots to verifiably solve long-horizon tasks with complex spatiotemporal constraints (Figure 1). Our approach has a modular structure consisting of a Language Instruction Module, a Spatial Grounding Module and a Task and Motion Planning Module. The first two modules utilize pretrained foundation models, while the third involves a task and motion planner for sequencing task subgoals and generating constraint-satisfying motion plans.
5.1 Language Instruction Module
Figure 2: An instruction is first translated into a conventional LTL formula $\phi_{l}$ that loosely captures the desired temporal occurrence of referent objects, then into our LTL syntax $\varphi_{l}$ with predicate functions that temporally chain required robot skills parameterized by composible referent descriptors.
In this module, we leverage a large language model $\psi$ to translate a natural language instruction $l$ into a linear temporal logic specification $\varphi_{l}$ with a novel composable syntax that enables referent disambiguation. We achieve this with a two-stage in-context learning strategy, the first stage translates $l$ into a conventional LTL formula $\phi_{l}$ where propositions refer to open-vocabulary objects. The second stage then takes as input $l$ and $\phi_{l}$ to generate a new formula $\varphi_{l}$, with predicate functions that correspond to parameterized robot skills. We define three predicate functions: near, pick and release for the primitive navigation and manipulation skills (picking and placing) required for multi-object goal navigation and mobile pick-and-place. Predicate functions in $\varphi_{l}$ are parameterized by Composable Referent Descriptors (CRDs), our novel propositional expressions that represent specific referent instances by chaining comparators that encode descriptive spatial information. For example the statement “the yellow cabinet above the fridge that is next to the stove”, can be represented with the composible referent descriptor below:
@@ 5.1 Language Instruction Module
which specifies that there is a fridge that is next to a stove, and the specific yellow cabinet we desire is above that fridge. CRDs are constructed from a set of 3D spatial comparators [29] defined in our prompting strategy. It is important to note that unlike recent works [9, 30], instructions do not require any specific phrasing syntax nor specific keywords and may feature arbitrary complexity and ambiguity. The LLM $\psi$ directly samples the entire LTL formula $\varphi_{l}$ with predicate functions parameterized by composable referent descriptors with appropriate spatial comparators based on the context of the instruction. Figure 2 presents an example highlighting the result of our two-stage prompting strategy. See appendix section A3 for further details.
@@
LLM Verification: Verifying the LTL formula $\varphi_{l}$ sampled from the LLM is vital as errors in referent extraction and temporal task structure implied by the sampled formula adversely affect accurate instruction following. To that end, the symbol verification node highlighted in Figure 1 leverages the unique properties of LTL to provide high-level human-in-the-loop verification for the extracted instruction referents and temporal task structure. Recent work [31] has addressed providing ISO 61508 [32] standard safety guarantees in robot task execution by translating descriptions of safety constraints from natural language to LTL formulas which are verified to be correct by human experts and used to enforce robot behavior. Similarly, we rely on human checking to verify that the translated formula $\varphi_{l}$ is correct, the symbol verification node implements an interactive dialog system that presents users with the extracted referent CRDs in the LTL formula as well as the implied task structure and reprompts the LLM based on corrective statements of the user to obtain new formulas. Unlike [31] we improve readability and relax the need for experts by directly translating the task structure––encoded in the sequence of state-machine transition expressions that must hold to progressively solve the task––back into English statements via a simple deterministic translation scheme. In our experiments (Table 1 & 2), we find that even without this human verification and reprompting step, the initial formulas sampled by our language understanding module impressively encodes the correct referents and implied temporal task structure. See appendix section A3.1 for further details.
5.2 Spatial Grounding Module

### v3

@@ Abstract
When instructing robots, users want to flexibly express constraints, refer to arbitrary landmarks, and verify robot behavior, while robots must disambiguate instructions into specifications and ground instruction referents in the real world. To address this problem, we propose Language Instruction grounding for Motion Planning (LIMP), an approach that enables robots to verifiably follow complex, open-ended instructions in real-world environments without prebuilt semantic maps. LIMP constructs a symbolic instruction representation that reveals the robot’s alignment with an instructor’s intended motives and affords the synthesis of correct-by-construction robot behaviors. We conduct a large-scale evaluation of LIMP on 150 instructions across five real-world environments, demonstrating its versatility and ease of deployment in diverse, unstructured domains. LIMP performs comparably to state-of-the-art baselines on standard open-vocabulary tasks and additionally achieves a 79% success rate on complex spatiotemporal instructions, significantly outperforming baselines that only reach 38%.  [^1] See supplementary materials and demo videos at robotlimp.github.io
I Introduction
@@
Robots need a rich understanding of natural language to be instructable by non-experts in unstructured environments. People, on the other hand, need to be able to verify that a robot has understood a given instruction and will act appropriately. Achieving these objectives, however, is challenging as natural language instructions often feature ambiguous phrasing, intricate spatiotemporal constraints, and unique referents. To illustrate, consider the instruction shown in Figure 1: “Bring the green plush toy to the whiteboard in front of it, watch out for the robot in front of the toy”. Solving such a task requires a robot to ground open-vocabulary referents, follow temporal constraints, and disambiguate objects using spatial descriptions. Foundation models [1, 2] offer a path to achieving such complex long-horizon goals; however, existing approaches for robot instruction following have largely focused on navigation [3, 4, 5, 6, 7]. These methods, broadly classified under object goal navigation [8], enable navigation to instances of an object category but are limited in their ability to localize spatial references and disambiguate object instances based on descriptive language. Other works [9, 10, 11] extend instruction following to mobile manipulation but are limited to tasks with simple temporal constraints expressed in unambiguous language. Moreover, existing efforts typically rely on Large Language Models (LLMs) as complete planners, bypassing intermediate symbolic representations that could provide verification of correctness before execution. Alternative approaches leveraging code-writing LLMs [5, 6, 12] are susceptible to errors in generated code, which may lead to unsafe robot behaviors. Mapping natural language to specification languages like temporal logic [13] provides a robust framework for language disambiguation, handling complex temporal constraints, and behavior verification. However, prior works along this line require prebuilt semantic maps with discrete sets of prespecified referents/landmarks from which instructions can be constructed [7, 14, 15].
Figure 1: Our approach executing the instruction “Bring the green plush toy to the whiteboard in front of it, watch out for the robot in front of the toy”. The robot dynamically detects and grounds open-vocabulary referents with spatial constraints to construct an instruction-specific semantic map, then synthesizes a task and motion plan to solve the task. In this example, the robot navigates from its start location (yellow, A), to the green plush toy (green, B), executes a pick skill then navigates to the whiteboard (blue, C), and executes a place skill. Note that the robot has no prior semantic knowledge of the environment.
We propose Language Instruction grounding for Motion Planning (LIMP), a method that leverages foundation models and temporal logics to dynamically generate instruction-conditioned semantic maps that enable robots to construct verifiable controllers for following navigation and mobile manipulation instructions with open vocabulary referents and complex spatiotemporal constraints. In a novel environment, LIMP constructs a 3D map via SLAM, then uses LLMs to translate complex natural language instructions into temporal logic specifications with a novel composable syntax for referent disambiguation. Instruction referents are detected and grounded using vision-language models (VLMs) and spatial reasoning. Finally, a task and motion plan is synthesized to guide the robot through the required subgoals, as shown in Figure 1. In summary, we make the following contributions: (1) A modular framework that translates expressive natural language instructions into temporal logic, grounds instruction referents, and executes commands via Task and Motion Planning (TAMP). (2) A spatial grounding method for detecting and localizing open vocabulary objects with spatial constraints in 3D metric maps. (3) A TAMP algorithm that localizes regions of interest (goal/avoidance zones) and synthesizes constraint-satisfying motion plans for long-horizon tasks.
II Background and Related Works
We briefly highlight the most relevant works in visual scene understanding [10], natural language instruction following [7, 16], and task and motion planning [17], and provide a comprehensive review in our supplementary materials. NLMap [10] grounds open-vocabulary language queries to spatial locations using pre-trained VLMs. While effective for describing individual objects, it cannot handle instructions involving complex constraints between multiple objects due to the lack of object relationship modeling. LIMP addresses this with a novel spatial grounding module that resolves spatial relationships and leverages task and motion planners to satisfy these constraints. Lang2LTL [7] is a multi-stage, LLM-based approach that uses entity extraction and replacement to translate language instructions into temporal logic. Its extension [16] incorporates VLMs and semantic information (via text embeddings) to ground referents. These works require prebuilt semantic maps/databases describing landmarks to ground symbols, whereas our approach dynamically generates landmarks based on open-vocabulary instructions. Action-Oriented Semantic Maps (AOSMs) [17] augment semantic maps with models indicating where robots can perform manipulation skills, integrating with TAMP solvers for mobile manipulation. LIMP similarly provides a TAMP-compatible spatial representation but supports open-vocabulary tasks, whereas AOSMs remain constrained to a fixed set of goals once generated.
II-A Linear Temporal Logic
LIMP translates natural language instructions into temporal logic specifications for verifiable task and motion planning. While compatible with various specification languages and planning frameworks, we choose Linear Temporal Logic (LTL) [18] for its proven expressivity in representing complex robot mission requirements [19]. LTL defines temporal properties using atomic propositions, logical operators—negation ($\neg$), conjunction ($\land$), disjunction ($\lor$), implication ($\rightarrow$)—and temporal operators: next ($\mathcal{X}$), until ($\mathcal{U}$), globally ($\mathcal{G}$), and finally ($\mathcal{F}$). Despite its expressivity, LTL has been underutilized due to the expert knowledge required to construct specifications, however recent works have seen significant success directly translating natural language into LTL [7, 20, 14, 21, 22, 23].
Behavior Verification: Expressing instructions as temporal logic specifications allows us to verify the correctness of generated plans a priori. However, instead of explicit verification methods such as model checking, we leverage insights from prior works [24] and directly use specifications to synthesize plans that are correct-by-construction [25, 26].
III Problem Definition
Given a natural language instruction $l$, our goal is to synthesize and sequence navigation and manipulation behaviors to produce a policy that satisfies the temporal and spatial constraints in $l$. Spatial constraints determine task success based on the sequence of robot poses traversed during execution; temporal constraints determine the sequencing of these spatial constraints as a function of task progression. We assume a robot with an RGB-D camera has already navigated a space, capturing images and camera poses. From this data, we build a metric map $m$ (e.g., point cloud, 3D voxel grid) of the environment, defining the space of possible $SE(3)$ poses $P$ and enabling robot localization (i.e., estimating $p_{\text{robot}}\in P$). Unlike previous work leveraging temporal logic [7], we do not assume access to a semantic map with prespecified object locations or predicates. Instead, we leverage two foundation models: a task-agnostic vision-language model $\sigma$ that, given an image and text, provides bounding boxes or segmentations based on the text; and an auto-regressive large language model $\psi$ that samples likely language tokens based on a history of tokens.
Navigation: Navigation is formalized as an object-goal oriented continuous path planning problem, where the goal is to generate paths to a goal pose set $P_{goals}\subset P$ while staying in feasible regions ($P_{feasible}\subset P$) and avoiding infeasible regions ($P_{infeasible}=P_{feasible}^{C}$). Infeasible regions include environment obstacles as well as dynamically determined semantic regions that violate constraints in the instruction $l$.
Manipulation: We formalize manipulation behaviours as options [27] parameterized by objects. Consider an object parameter $\theta$ that parameterizes an option $O_{\theta}=(I_{\theta},\pi_{\theta},\beta_{\theta})$, the initiation set, policy, and termination condition are functions of both the robot pose $P$ and $\theta$. The initiation set $I_{\theta}$ denotes the global reference frame robot positions and object-centric attributes––such as object size––that determine if the option policy $\pi_{\theta}$ can be executed on the object $\theta$. To execute a manipulation skill on an object, an object-goal navigation behavior must first be executed to bring the robot into proximity with the object. We assume access to a library of these manipulation skills and demonstrate our approach on multi-object goal navigation and open-vocabulary mobile pick-and-place [9, 28].
IV Language Instruction Grounding for Motion Planning
Figure 2: [A] LIMP translates natural language instructions into temporal logic expressions, where open-vocabulary referents are applied to predicates that correspond to robot skills––note the context-aware resolution of the phrase “blue one” to the referent “blue_sofa”. [B] Vision-language models detect referents, while spatial reasoning disambiguates referent instances to generate a 3D semantic map that localizes instruction-specific referents. [C] Finally, the temporal logic expression is compiled into a finite-state automaton, which a task and motion planner uses with dynamically-generated task progression semantic maps to progressively identify goals and constraints in the environment, and generate a plan that satisfies the high-level task specification.
LIMP interprets expressive natural language instructions to generate instruction-conditioned semantic maps, enabling robots to verifiably solve long-horizon tasks with complex spatiotemporal constraints (Figure 2). We briefly describe our modular approach in this section and present comprehensive implementations details in our supplementary materials.
IV-A Language Instruction Module
In this module, we leverage a large language model $\psi$ to translate a natural language instruction $l$ into a linear temporal logic specification $\varphi_{l}$ with a novel composable syntax for referent disambiguation. We achieve this through a two-stage in-context learning strategy. The first stage prompts $\psi$ to translate $l$ into a conventional LTL formula $\phi_{l}$ where propositions refer to open-vocabulary objects. The second stage takes $l$ and $\phi_{l}$ as input and prompts $\psi$ to generate a new formula $\varphi_{l}$ with predicate functions corresponding to parameterized robot skills.
We define three predicate functions—near, pick, and release—for the primitive navigation and manipulation skills required for multi-object goal navigation and mobile pick-and-place. Predicate functions in $\varphi_{l}$ are parameterized by Composable Referent Descriptors (CRDs), our novel propositional expressions representing specific referent instances by chaining comparators that encode descriptive spatial information. For example, the instruction “the yellow cabinet above the fridge that is next to the stove” can be represented with the CRD:
[…]
This specifies that there is a fridge next to a stove, and the desired yellow cabinet is above that fridge. CRDs are constructed from a set of 3D spatial comparators [29] defined in our prompting strategy.
@@ IV-A Language Instruction Module
Unlike recent works [9, 11], our approach does not require specific phrasing or keywords and can handle instructions with arbitrary complexity and ambiguity. The LLM $\psi$ directly samples the entire LTL formula $\varphi_{l}$ with predicate functions parameterized by CRDs using appropriate spatial comparators based on the instruction’s context. Figure 3 illustrates the result of our two-stage prompting strategy.
Figure 3: An instruction is first translated into a conventional LTL formula $\phi_{l}$ that loosely captures the desired temporal occurrence of referent objects, then into our LTL syntax $\varphi_{l}$ with predicate functions that temporally chain required robot skills parameterized by composable referent descriptors.
@@
LLM Verification: Verifying the LTL formula $\varphi_{l}$ sampled from the LLM is crucial as errors in referent extraction and temporal task structure affects instruction following accuracy. Our symbol verification node (Figure 2) leverages LTL properties to provide high-level human-in-the-loop verification of extracted instruction referents and temporal task structure. Recent work [30] provides ISO 61508 [31] safety guarantees in robot task execution by translating safety constraints from natural language to LTL formulas, which are verified by human experts and used to enforce robot behavior. Similarly, we rely on human verification to ensure the translated formula $\varphi_{l}$ is correct. Our symbol verification node implements an interactive dialog system that presents users with the extracted referent CRDs and implied task structure, and reprompts the LLM based on user corrections to obtain new formulas. Unlike prior work [30], we eliminate the need for experts by directly translating the task structure—encoded in the LTL formulas’s equivalent automaton—back into English statements via a simple deterministic translation scheme. In our experiments (Tables I and II), we find that even without human verification and reprompting, the initial formulas sampled by our language understanding module impressively encode the correct referents and temporal task structure.
IV-B Spatial Grounding Module